# A Unified Model for Multi-Task Drone Routing in Post-Disaster Road Assessment

### v1

@@ 7.3 Sensitivity analysis
To evaluate the robustness and scalability of the proposed UM, this section conducts sensitivity analyses across three critical dimensions: drone fleet sizes, maximum allowable assessment time constraints, and network scales. These analyses assess performance stability when key parameters deviate from training conditions, demonstrating the model’s applicability to diverse post-disaster scenarios. All sensitivity tests are conducted on the PDRA-OR-TW variant, which simultaneously incorporates the open route and time window constraints, thereby representing a complex and practically relevant scenario. The results yield three key insights:
@@
    Table 6 shows performance sensitivity to varying drone fleet sizes (4, 5, 6, and 7 drones) on a fixed 400-node network with $p_{max}=45$ minutes. This analysis is particularly critical for disaster response scenarios, where drone availability often fluctuates due to uncertainties. Table 6 reveals that UM maintains consistent superiority across varying drone fleet sizes, with performance gaps ranging from 1.20% to 10.34% over the single-task AEDM baselines. Gurobi fails to produce feasible solutions within 60 seconds for all fleet configurations, highlighting UM’s advantage in time-critical scenarios. The computational time scales linearly with fleet size (2–4 seconds), demonstrating efficient parallelization capabilities essential for large-scale disaster response operations.
@@
    Table 7 shows performance under varying maximum allowable assessment times (20, 30, 40, and 50 minutes) with fixed network scale (400 nodes) and drone fleet size (4). This analysis simulates the time-critical nature of disaster response operations where assessment windows may vary based on emergency severity and resource availability. The results demonstrate UM’s robustness under temporal constraints, maintaining superior performance across all maximum allowable assessment time constraints with gaps ranging from -0.20% to 16.07% over single-task AEDM baselines. Gurobi consistently fails at 60-second limits across all time windows, requiring 10–30 minutes to achieve suboptimal solutions.
@@
    Table 8 shows UM’s scalability across varying network scales (400, 600, 800, and 1,000 nodes) while maintaining fixed assessment time ($p_{max}=45$ minutes) and drone fleet size (4). The results reveal strong scalability characteristics, with UM preserving solution quality superiority across all tested network scales. Performance gaps over the single-task AEDM baselines range from 0.17% to 11.65%, while its computational time remains within real-time limits (2–10 seconds). In contrast, Gurobi fails completely on networks larger than 400 nodes, even with extended time limits up to 30×60 seconds, underscoring UM’s practical advantage for large-scale PDRA applications where traditional optimization methods become computationally intractable.
Drone fleet sizes
Method 4 5 6 7
Gurobi 100% 60 100% 60 100% 60 100% 60
15.20 65.54% 10×60 15.28 69.93% 10×60 18.92 66.13% 10×60 21.56 63.66% 10×60
17.93 59.35% 20×60 24.50 51.79% 20×60 28.65 48.71% 20×60 30.89 47.93% 20×60
25.29 42.67% 30×60 25.14 50.53% 30×60 32.76 41.35% 30×60 34.09 42.54% 30×60
AEDM-Basic 40.37 8.48% 2 46.98 7.56% 2 50.28 9.99% 3 53.91 9.14% 4
AEDM-OR 39.55 10.34% 2 46.46 8.58% 2 50.20 10.13% 3 53.41 9.98% 4
AEDM-TW 42.00 4.78% 2 48.84 3.90% 2 52.37 6.25% 3 56.07 5.49% 4
AEDM-OR-TW 43.33 1.77% 2 49.00 3.58% 2 55.19 1.20% 3 58.57 1.28% 4
UM 44.11 0.00% 2 50.82 0.00% 2 55.86 0.00% 3 59.33 0.00% 4
@@
The sensitivity analyses collectively validate UM’s robustness across critical operational parameters, confirming its suitability for deployment in diverse post-disaster environments where parameter variations are inevitable. The consistent performance superiority and computational efficiency across all tested dimensions demonstrate the model’s practical reliability for time-sensitive emergency response operations where both solution quality and computational speed are important.
7.4 Real-world application
To further validate the practical applicability of the proposed UM in real-world post-disaster scenarios, experiments are conducted on the publicly available Anaheim transportation network (as shown in Figure 6), with data sourced from https://github.com/bstabler/TransportationNetworks. This network simulates the complexity of a metropolitan road system, which is typical of large-scale PDRAs. The original network (416 nodes and 914 links) is transformed into a 1330-node structure (including 914 artificial nodes for road links) using the transformation method detailed in Section 3.2. The experimental setup includes 7 drones, and the maximum assessment time $p_{max}$ is set to 45 minutes. Three benchmark approaches are used: UM, single-task AEDMs (Gong et al. 2025), and Gurobi. In the experiments, UM solves these 4 PDRA variants in a unified manner, while the four single-task AEDMs solve all variants (with reported results being their average values); additionally, Gurobi’s time limit is extended to 10,000 seconds to account for the increased problem complexity. Results in Figure 7 show that UM consistently outperforms the other two benchmarks: compared with the single-task AEDMs, it achieves an average performance improvement of 7.9% (ranging from 7.2% to 8.4%) by leveraging multi-task knowledge sharing, and relative to Gurobi, it yields a 25.5% improvement (ranging from 23.6% to 27.8%). In terms of efficiency, UM only takes approximately 10 seconds to solve the problem, while the single-task AEDMs require variant-specific deployment; Gurobi’s 10,000-second runtime, meanwhile, is entirely impractical for real-world use. These findings confirm UM’s scalability and robustness under realistic network complexity, as well as its flexibility for rapid deployment in time-critical PDRA applications.
Figure 6: Anaheim Network
Figure 7: Performance Gap Comparison of UM, AEDM, and Gurobi Across Four PDRA Variants on the Anaheim Network
@@ 7.5 Finetuning
To evaluate the proposed lightweight finetuning mechanism for incorporating previously unseen attributes, experiments are conducted on 400-node networks with 2 depots, each deploying 3 drones under a maximum assessment time constraint of $p_{max}=45$ minutes. The MD attribute, deliberately excluded from initial UM training, serves as the representative unseen attribute to assess adaptation capabilities. The finetuning mechanism leverages efficient adapter layers through parameter augmentation with zero-initialized entries, requiring only minimal modifications to the input embedding and decoder’s context embedding while preserving the valuable multi-task knowledge encoded in the original parameters. Four approaches are systematically compared: (i) Gurobi, representing traditional optimization; (ii) single-task AEDM variants (AEDM-MD, AEDM-OR-MD, AEDM-TW-MD, and AEDM-OR-TW-MD), each trained from scratch for 200 epochs; (iii) UM-Zero-Shot, where the pre-trained UM is directly applied to MD problems without adaptation; and (iv) UM-10-Epochs, which applies the lightweight adapter mechanism with only 10 epochs of finetuning. The results are summarized in Table 9.
@@
Value Gap Time (s) Value Gap Time (s) Value Gap Time Value Gap Time (s)
@@
Compared with Gurobi, UM-10-Epochs shows advantages in both solution quality and efficiency. Gurobi often fails to produce feasible solutions within the 60-second limit and requires 10–30 minutes to obtain suboptimal outcomes, whereas UM-10-Epochs consistently generates high-quality solutions in just a few seconds. This indicates that the adapter mechanism effectively integrates the unseen MD attribute without compromising real-time performance, which is essential in post-disaster contexts. When compared with single-task AEDM variants, UM-10-Epochs achieves higher solution quality despite AEDMs undergoing full training on each MD-integrated problem. This highlights that multi-task knowledge transfer, coupled with lightweight finetuning, allows UM to learn new attribute-specific knowledge at only a fraction of the training cost, avoiding the inefficiency of maintaining multiple specialized models. Relative to UM-Zero-Shot, UM-10-Epochs eliminates the moderate performance gap that arises when the MD attribute is directly introduced without adaptation. The short 10-epoch finetuning is sufficient to fully close this gap, yielding optimal performance across all MD-integrated PDRA variants. Overall, these results confirm that even minimal finetuning enables UM to reconcile pre-trained knowledge with new operational constraints, validating the effectiveness of the adapter mechanism. UM thus proves particularly valuable for rapid adaptation in evolving disaster response scenarios, where new operational requirements may emerge with limited preparation time.
@@ 8 Conclusion
This study addresses the critical challenge of efficient drone routing for PDRA, where existing methods suffer from computational inefficiencies and lack adaptability to diverse operational scenarios. Traditional approaches, including exact optimization methods and single-task deep learning models, require separate solutions for each PDRA variant, resulting in substantial computational overhead and deployment complexity. We propose a UM that consolidates all PDRA variants within a single model through multi-task learning. UM employs a modern transformer architecture incorporating RMS normalization, pre-normalization configuration, FA, and SGLUFFN to enhance computational efficiency and solution quality. The unified approach eliminates the need for variant-specific models while leveraging shared knowledge across different operational constraints. Experimental evaluation demonstrates substantial improvements over existing approaches. UM achieves an 8-fold reduction in training time (24 vs. 192 hours) and model parameters (1.3M vs. 10.4M) compared to training separate models for each variant. Solution quality consistently surpasses single-task methods by 6–14% and traditional optimization approaches by 24–82%, while maintaining real-time performance (1–10 seconds) regardless of network scale or operational complexity. Sensitivity analyses confirm robust performance across varying drone fleet sizes, assessment time constraints, and network scales up to 1,000 nodes. The lightweight adapter mechanism enables efficient incorporation of previously unseen attributes through minimal parameter augmentation and brief finetuning, demonstrating practical adaptability to evolving disaster response requirements.
Future research can proceed in several directions. First, inspired by the scaling laws observed in LLM, where performance improves as parameter size grows (Achiam et al. 2023, Touvron et al. 2023, Guo et al. 2025), we plan to investigate whether increasing the parameter scale of UM (from millions to potentially billions) can yield further gains in solution quality, generalization, and adaptability. Second, beyond PDRA, the unified multi-task learning framework could be applied to broader classes of neural combinatorial optimization problems in emergency logistics, transportation planning, and other time-critical domains, contributing to the development of general-purpose decision-making models for complex real-world systems.
A. Model component details
@@
RMS normalization, originally introduced by Zhang and Sennrich (2019), provides improved training stability and reduced computational overhead compared to traditional instance normalization. In our encoder architecture, RMS normalization replaces instance normalization as described in Equation (3). For a given input vector $\mathbf{x}\in\mathbb{R}^{d}$, where $d$ represents the embedding dimension, RMS normalization is computed as:
@@
FA optimizes memory usage and computational efficiency in attention mechanisms without sacrificing numerical accuracy (Dao et al. 2022, Dao 2023), as referenced in Equations (4) and (7). The standard MHA computation is given by:
@@ B. Ablation study
Performance results measured by optimality gap relative to the full UM reveal significant component contributions. Figure 8 shows the ablation results across different problem instances. The Base Group with traditional components achieves the poorest performance, demonstrating substantial optimality gaps. Individual component removals show varying impacts: removing RMS normalization (-RMS) causes moderate performance degradation, while removing pre-normalization (-PreNorm) leads to more significant deterioration. The SGLUFFN activation function proves particularly critical, as its removal (-SGLUFFN) results in notable performance drops. FA removal (-FA) shows the least impact among individual components, though still measurable. These results validate our architectural choices and demonstrate that each modern component contributes meaningfully to the overall model performance.
Figure 8: Ablation Study Results Showing Optimality Gaps for Different Architectural Variants
C. Mathematical formulation
@@
Let $x_{ij}^{k}$ be the binary decision variable indicating whether drone $k$ traverses link $(i,j)$ for assessment, where $x_{ij}^{k}=1$ if drone $k$ flies the link, and $x_{ij}^{k}=0$ otherwise. The PDRA-Basic model is formulated as (Gong et al. 2025):
$\max\quad$ $\sum_{k\in K}\sum_{p\in P}c_{p}\left(\sum_{j:(p,j)\in\bar{A}}x_{pj}^{k}\right)$ (C1)
s.t. $\sum_{k\in K}\sum_{j:(p,j)\in\bar{A}}x_{pj}^{k}\leq 1\quad\forall p\in P$ (C2)
@@
$u_{i}^{k}+1-u_{j}^{k}\leq M(1-x_{ij}^{k})\quad\forall(i,j)\in\bar{A},\forall k\in K$ (C7)
@@
$u_{i}^{k}\in\{1,2,\ldots,|\bar{N}|\}\quad\forall i\in\bar{N},\forall k\in K$ (C9)
@@
  • Equation (C1): Maximize total information value collected from artificial nodes $P$
@@
  • Equation (C7): Miller-Tucker-Zemlin subtour elimination constraint with $M=|\bar{N}|$
@@
  • $\delta_{o}$: Capacity (maximum drones) at depot $o$
@@
where Equation (C15) assigns each drone to at most one depot, Equation (C16) links depot assignment to path decisions, Equation (C17) enforces depot capacity, Equation (C18) ensures return to assigned depot, and Equation (C19) defines binary depot assignment variables.

### v2

@@ 7.3 Sensitivity analysis
To evaluate the robustness and scalability of the proposed UM, this section conducts sensitivity analyses across three critical dimensions: (i) maximum allowable assessment time, (ii) drone fleet size, and (iii) network scale. These analyses assess performance stability when key parameters deviate from training conditions, demonstrating the model’s applicability to diverse post-disaster scenarios. All sensitivity tests are conducted on the PDRA-OR-TW variant, which simultaneously incorporates the open route and time window constraints, thereby representing a complex and practically relevant scenario. The results yield three key insights:
@@
    Table 6 reports performance under varying assessment time limits (20, 30, 40, and 50 minutes) on a fixed 400-node network with four drones. This experiment reflects the time-critical nature of PDRA missions, where available assessment windows may vary by disaster severity or resource constraints. UM maintains superior solution quality across all tested settings, with performance gaps over the best single-task AEDM ranging from approximately $0.20\%$ to $16.07\%$. Gurobi fails to produce feasible solutions within a 60-second limit across all time windows, requiring full 30-minute runs to return suboptimal results. The GC and GC+LS heuristics remain consistently inferior to UM, although GC+LS achieves noticeable improvements over GC. These findings confirm UM’s robustness to temporal variability in PDRA operations.
@@
    Table 7 examines scenarios with 4, 6, 8, and 10 drones on a fixed 400-node network and $p_{\max}=45$ minutes, simulating practical fluctuations in drone availability during disaster response. UM consistently outperforms all baselines, with performance gaps of $1.20\%$$12.79\%$ relative to the best AEDM model. Gurobi again fails to produce feasible solutions within the 60-second limit for all fleet sizes, highlighting UM’s advantage in time-critical PDRA settings. The GC and GC+LS heuristics show predictable degradation as complexity increases, though GC+LS remains noticeably stronger than GC. UM’s computational time scales linearly with the number of drones (approximately 2–10 seconds), illustrating efficient parallelization capability for multi-drone coordination.
@@
    Table 8 evaluates scalability on networks of 400, 600, 800, and 1,000 nodes while keeping $p_{\max}=45$ minutes and the drone fleet size fixed at four drones. UM demonstrates strong scalability, preserving solution-quality superiority across all scales, with performance gaps of $0.17\%$$11.65\%$ over the best AEDM. The GC and GC+LS heuristics show increasing deterioration as network size grows, reflecting their structural limitations under large-scale PDRA settings. Gurobi becomes completely intractable beyond 400 nodes, failing even with an extended 30-minute limit. Despite increased problem size, UM maintains rapid computational performance (2–10 seconds), underscoring its practical viability for large-scale PDRA applications.
@@
Drone fleet sizes
Method 4 6 8 10
Gurobi 100% 60 100% 60 100% 60 100% 60
25.29 42.67% 30×60 32.76 41.35% 30×60 38.50 40.28% 30×60 42.80 39.45% 30×60
GC-OR-TW 15.20 65.54% 24 21.35 61.78% 57 25.80 59.90% 95 29.50 58.35% 138
GC+LS-OR-TW 23.12 47.60% 620 38.45 31.17% 1501 48.20 25.24% 2485 55.80 21.19% 3580
AEDM-Basic 40.37 8.48% 2 50.28 9.99% 3 56.80 11.83% 6 62.15 12.16% 10
AEDM-OR 39.55 10.34% 2 50.20 10.13% 3 56.50 12.31% 6 61.70 12.79% 10
AEDM-TW 42.00 4.78% 2 52.37 6.25% 3 58.95 8.47% 6 64.30 9.14% 10
AEDM-OR-TW 43.33 1.77% 2 55.19 1.20% 3 61.85 3.98% 6 67.45 4.66% 10
UM 44.11 0.00% 2 55.86 0.00% 3 64.42 0.00% 6 70.78 0.00% 10
@@
The sensitivity analyses collectively validate UM’s robustness across critical operational parameters, confirming its suitability for deployment in diverse post-disaster environments where parameter variations are inevitable. Across all tested dimensions of temporal constraints, fleet sizes, and network scales, UM demonstrates consistent performance superiority and computational efficiency, confirming the model’s practical reliability for time-sensitive emergency response operations where both solution quality and computational speed are paramount. Furthermore, the comparisons with both heuristic methods and exact optimization highlight the fundamental limitations of traditional approaches: heuristics require extensive domain expertise and variant-specific design yet achieve suboptimal solutions, while exact methods suffer from computational intractability on realistic problem scales. UM successfully overcomes both limitations through learned multi-task representations that generalize effectively across diverse operational scenarios.
@@ 7.4 Finetuning
To evaluate the proposed lightweight finetuning mechanism for incorporating previously unseen attributes, experiments are conducted on 400-node networks with 2 depots, each deploying 3 drones under a maximum assessment time constraint of $p_{max}=45$ minutes. The MD attribute, deliberately excluded from initial UM training, serves as the representative unseen attribute to assess adaptation capabilities. The finetuning mechanism leverages efficient adapter layers through parameter augmentation with zero-initialized entries, requiring only minimal modifications to the input embedding and decoder’s context embedding while preserving the valuable multi-task knowledge encoded in the original parameters. Five approaches are compared: (i) Gurobi, representing exact optimization; (ii) GC and GC+LS heuristics, representing heuristic optimization with their variant-specific implementations for MD-integrated problems (GC-Variant and GC+LS-Variant); (iii) single-task AEDM variants (AEDM-MD, AEDM-OR-MD, AEDM-TW-MD, and AEDM-OR-TW-MD), each trained from scratch for 200 epochs; (iv) UM-Zero-Shot, where the pre-trained UM is directly applied to MD problems without adaptation; and (v) UM-10-Epochs, which applies the lightweight adapter mechanism with only 10 epochs of finetuning. The results are summarized in Table 9.
Compared with Gurobi, UM-10-Epochs shows advantages in both solution quality and efficiency. Gurobi often fails to produce feasible solutions within the 60-second limit and requires full 30 minutes to obtain suboptimal outcomes, whereas UM-10-Epochs consistently generates high-quality solutions in just a few seconds. For heuristic methods, GC+LS achieves better performance than GC alone through local search improvement, but both remain substantially inferior to UM-10-Epochs. When compared with single-task AEDM variants, UM-10-Epochs achieves higher solution quality despite AEDMs undergoing full training on each MD-integrated problem. This highlights that multi-task knowledge transfer, coupled with lightweight finetuning, allows UM to learn new attribute-specific knowledge at only a fraction of the training cost, avoiding the inefficiency of maintaining multiple specialized models. Relative to UM-Zero-Shot, UM-10-Epochs eliminates the moderate performance gap that arises when the MD attribute is directly introduced without adaptation. The short 10-epoch finetuning is sufficient to fully close this gap, yielding better performance across all MD-integrated PDRA variants. Overall, these results confirm that even minimal finetuning enables UM to reconcile pre-trained knowledge with new operational constraints, validating the effectiveness of the adapter mechanism. UM thus proves particularly valuable for rapid adaptation in evolving disaster response scenarios, where new operational requirements may emerge with limited preparation time.
@@
7.5 Real-world application
Figure 6: Real-World Anaheim Transportation Network
To further validate the practical applicability of the proposed UM in real-world post-disaster scenarios, experiments are conducted on the publicly available Anaheim transportation network (as shown in Figure 6), with data sourced from https://github.com/bstabler/TransportationNetworks. This network simulates the complexity of a road system typical of large-scale PDRAs. The original network (416 nodes and 914 links) is transformed into a 1330-node structure (including 914 artificial nodes for road links) using the transformation method detailed in Section 3.2. The experimental setup includes 10 drones, and the maximum assessment time $p_{\max}$ is set to 45 minutes. Four benchmark approaches are evaluated: (i) UM, solving all 8 PDRA variants in a unified manner; (ii) single-task AEDMs, with 8 separate models trained for each variant; (iii) GC and GC+LS heuristics, with 8 variant-specific implementations; and (iv) Gurobi, with the time limit extended to 10,000 seconds to account for increased problem complexity.
Figure 7: Solution Quality and Computational Efficiency on Anaheim Network
Figure 7 presents the comparative results. Two key observations emerge from the real-world network evaluation. As shown in Figure 7 (a), UM consistently achieves superior solution quality compared to all baseline methods on the large-scale Anaheim network. UM outperforms the best single-task AEDM, demonstrating that the unified multi-task learning approach maintains its advantage even on realistic network topologies. Compared to heuristic methods, UM shows substantial improvements over both GC+LS and GC. These results confirm that hand-crafted heuristics, despite incorporating local search improvements, remain fundamentally limited by their reliance on predefined rules that fail to capture the complex patterns inherent in real-world road networks. Most notably, Gurobi exhibits the poorest performance among all methods. Even with extended time limits, Gurobi frequently fails to generate high-quality feasible solutions for the large-scale network, highlighting the computational intractability of exact optimization methods for real-world large-scale PDRA applications. Figure 7 (b) summarizes the average computational time of these methods for solving the 8 variants and illustrates the dramatic computational efficiency advantages of UM for real-world deployment. UM maintains rapid performance across all PDRA variants. In stark contrast, Gurobi requires extended time limits to return suboptimal solutions, making it impractical for time-critical disaster response scenarios where decisions must be made within minutes. Heuristic methods demonstrate intermediate performance: GC provides rapid construction but at the cost of poor solution quality, while GC+LS achieves moderate improvements through local search but requires substantially longer computation time. Notably, despite requiring separate models for each variant and higher total parameter counts, single-task AEDMs achieve comparable inference speed to UM due to the inherently rapid nature of neural network forward passes. However, UM’s unified architecture offers critical advantages in deployment flexibility and training efficiency. By requiring only one model with significantly fewer parameters and reduced training time, UM maintains equivalent rapid performance and superior solution quality compared to multiple specialized models. These findings confirm UM’s scalability and robustness under realistic network complexity. The model successfully handles the irregular topology, varied connectivity patterns, and large-scale structure of real-world transportation networks while maintaining both solution quality superiority and rapid computational performance. The results validate UM’s practical viability for rapid deployment in time-critical PDRA applications, where both high-quality routing decisions and rapid response times are essential for effective disaster response operations.
@@ 8 Conclusion
This study addresses the critical challenge of efficient drone routing for PDRA, where existing methods suffer from computational inefficiencies and lack adaptability to diverse operational scenarios. Traditional approaches, including exact optimization methods, heuristic optimization methods, and single-task deep learning models, require separate solutions for each PDRA variant, resulting in substantial computational overhead and deployment complexity. We propose a UM that consolidates all PDRA variants within a single model through multi-task learning. UM employs a modern transformer architecture incorporating RMS normalization, pre-normalization configuration, FA, and SGLUFFN to enhance computational efficiency and solution quality. The unified approach eliminates the need for variant-specific models while leveraging shared knowledge across different operational constraints. Experimental evaluation demonstrates substantial improvements over existing approaches. UM achieves an 8-fold reduction in training time (24 vs. 192 hours) and model parameters (1.3M vs. 10.4M) compared to training separate models for each variant. Solution quality consistently surpasses single-task DRL methods by 6–14%, heuristic algorithms by 22–42% and commercial solvers by 24–82%, while maintaining rapid performance (1–10 seconds) regardless of network scale or operational complexity. Sensitivity analyses confirm robust performance across varying drone fleet sizes, assessment time constraints, and network scales up to 1,000 nodes. The lightweight adapter mechanism enables efficient incorporation of previously unseen attributes through minimal parameter augmentation and brief finetuning, demonstrating practical adaptability to evolving disaster response requirements.
Future research can proceed in several directions. First, inspired by the scaling laws observed in LLM, where performance improves as parameter size grows (achiam2023gpt; touvron2023llama; guo2025deepseek), we plan to investigate whether increasing the parameter scale of UM (from millions to potentially billions) can yield further gains in solution quality, generalization, and adaptability. Second, beyond PDRA, the unified multi-task learning framework could be applied to broader classes of neural combinatorial optimization problems in emergency logistics, transportation planning, and other time-critical domains, contributing to the development of general-purpose decision-making models for complex real-world systems.
Appendix A Model component details
@@
RMS normalization, originally introduced by zhang2019root, provides improved training stability and reduced computational overhead compared to traditional instance normalization. In our encoder architecture, RMS normalization replaces instance normalization as described in Equation (3). For a given input vector $\mathbf{x}\in\mathbb{R}^{d}$, where $d$ represents the embedding dimension, RMS normalization is computed as:
@@
FA optimizes memory usage and computational efficiency in attention mechanisms without sacrificing numerical accuracy (dao2022flashattention; dao2023flashattention), as referenced in Equations (4) and (7). The standard MHA computation is given by:
@@ Appendix B Ablation study
Performance results measured by relative performance gap compared to the full UM reveal significant component contributions. Figure 8 shows the ablation results across different problem instances. The Base Group with traditional components achieves the poorest performance, demonstrating substantial performance gaps. Individual component removals show varying impacts: removing RMS normalization (-RMS) causes moderate performance degradation, while removing pre-normalization (-PreNorm) leads to more significant deterioration. The SGLUFFN activation function proves particularly critical, as its removal (-SGLUFFN) results in notable performance drops. FA removal (-FA) shows the least impact among individual components, though still measurable. These results validate our architectural choices and demonstrate that each modern component contributes meaningfully to the overall model performance.
Figure 8: Ablation Study Results Showing Performance Gaps for Different Architectural Variants
Appendix C Mathematical formulation
@@
Let $x_{ij}^{k}$ be the binary decision variable indicating whether drone $k$ traverses link $(i,j)$ for assessment, where $x_{ij}^{k}=1$ if drone $k$ flies the link, and $x_{ij}^{k}=0$ otherwise. Let $f_{ij}^{k}\geq 0$ be the continuous flow variable representing the flow carried by drone $k$ on link $(i,j)$. The PDRA-Basic model is formulated as (Gong2025Deep):
$\max\quad\sum_{k\in K}\sum_{p\in\mathcal{P}}c_{p}\left(\sum_{j:(p,j)\in\bar{A}}x_{pj}^{k}\right)$ (C1)
s.t. $\sum_{k\in K}\sum_{j:(p,j)\in\bar{A}}x_{pj}^{k}\leq 1\quad\forall p\in\mathcal{P}$ (C2)
@@
$\sum_{j:(o,j)\in\bar{A}}f_{oj}^{k}-\sum_{j:(j,o)\in\bar{A}}f_{jo}^{k}=\sum_{p\in\mathcal{P}}\sum_{j:(p,j)\in\bar{A}}x_{pj}^{k}\quad\forall k\in K$ (C8)
$\sum_{j:(j,i)\in\bar{A}}f_{ji}^{k}-\sum_{j:(i,j)\in\bar{A}}f_{ij}^{k}=\sum_{j:(i,j)\in\bar{A}}x_{ij}^{k}\quad\forall i\in\mathcal{P},\forall k\in K$ (C9)
$\sum_{j:(j,i)\in\bar{A}}f_{ji}^{k}-\sum_{j:(i,j)\in\bar{A}}f_{ij}^{k}=0\quad\forall i\in N\setminus\{o\},\forall k\in K$ (C10)
$f_{ij}^{k}\leq|\mathcal{P}|\cdot x_{ij}^{k}\quad\forall(i,j)\in\bar{A},\forall k\in K$ (C11)
@@
$f_{ij}^{k}\geq 0\quad\forall(i,j)\in\bar{A},\forall k\in K$ (C13)
@@
  • Equation (C1): Maximize total information value collected from artificial nodes $\mathcal{P}$
@@
  • Equation (C5): Prevent simultaneous bidirectional traversal of artificial nodes
@@
  • Equation (C8): Depot flow balance constraint with net outflow equal to number of visited artificial nodes
  • Equation (C9): Flow conservation at artificial nodes with unit consumption for assessment completion
  • Equation (C10): Flow conservation at original network nodes allowing revisits
  • Equation (C11): Flow capacity constraint linking flow variables to routing decisions
@@
  • $\delta_{o}$: Maximum number of drones available at depot $o$
@@
where Equation (C19) assigns each drone to at most one depot, Equation (C20) links depot assignment to departure decisions, Equation (C21) limits the number of drones at each depot, Equation (C22) ensures return to assigned depot, and Equation (C23) defines binary depot assignment variables.